# Failure of famous functional inequalities on Finsler manifolds: the influence of $S$-curvature

### v1

@@ 2.1. Elements from Finsler geometry
In this section we recall some definitions and properties from Finsler geometry; for details see e.g. Bao, Chern and Shen [2] and Shen [26, 27].
@@ 2.1.1. Finsler manifolds
A Minkowski norm on an $n$($\geq 2$)-dimensional vector space $V$ is a $C^{\infty}$-function $\phi:V\setminus\{0\}\rightarrow[0,+\infty)$ satisfying:
@@
which induces a Riemannian metric on $T_{x}M\backslash\{0\}$. The Finsler metric $F$ is (induced by) a Riemannian metric if and only if $g_{y}$ is independent of $y$, i.e., $g_{ij}(x,y)=g_{ij}(x)$. In the same way, $F$ is a Minkowski metric if $g_{y}$ is independent of $x$, i.e., $g_{ij}(x,y)=g_{ij}(y)$.
The reversibility of a set $U\subset M$, cf. Rademacher [24], is defined by
@@
Thus $\lambda_{F}(M)\geq 1,$ with equality if and only if $F$ is reversible, i.e., $F(x,y)=F(x,-y)$ for every $(x,y)\in TM$. Moreover, $x\mapsto\lambda_{F}(x)$ is continuous on $M$. In particular, $\lambda_{F}(M)$ is finite when $M$ is compact or $F$ is a Minkowski metric.
@@
where $g_{y}(y,\cdot):=g_{ij}(x,y)y^{i}{\operatorname{d}}x^{j}$. Given a smooth function $u:M\rightarrow\mathbb{R}$, the gradient of $u$ is
@@
In particular, for any smooth vector filed $X$, one has $\langle X,{\operatorname{d}}u\rangle=X(u)=g_{\nabla u}(\nabla u,X),$ where $\langle y,\xi\rangle:=\xi(y)$ denotes the canonical pairing between $T_{x}M$ and $T^{*}_{x}M$.
@@
$\langle y,\xi\rangle\leq F(x,y)F^{*}(x,\xi),\quad\text{ for any }y\in T_{x}M,\ \xi\in T^{*}_{x}M,$ (2.4)
with equality if and only if $\xi=\alpha\mathfrak{L}(x,y)$ for some $\alpha>0$. In addition, $F(x,y)=F^{*}(\mathfrak{L}(x,y))$ for any $y\in T_{x}M$. Furthermore, the reversibilities of $F$ and $F^{*}$ coincide, see Huang, Kristály and Zhao [10, Lemma 2.1], i.e.,
@@
The distance function $d_{F}:M\times M\rightarrow[0,\infty)$ is defined as $d_{F}(x_{1},x_{2}):=\inf L_{F}(c),$ where the infimum is taken over all locally Lipschitz curves $c:[0,1]\rightarrow M$ with $c(0)=x_{1}$ and $c(1)=x_{2}$. For any $x_{1},x_{2},x_{3}\in M$, one has that $d_{F}(x_{1},x_{2})\geq 0,$ with equality if and only if $x_{1}=x_{2}$, and $d_{F}(x_{1},x_{2})\leq d_{F}(x_{1},x_{3})+d_{F}(x_{3},x_{2})$. Usually $d_{F}(x_{1},x_{2})\neq d_{F}(x_{2},x_{1}),$ unless $F$ is reversible. In fact, one has for every $x_{1},x_{2}\in M$ that
@@
A Finsler manifold $(M,F)$ is called forward complete if every geodesic $t\mapsto\gamma(t)$, $t\in[0,1)$, can be extended to a geodesic defined on $t\in[0,+\infty)$; similarly, $(M,F)$ is called backward complete if every geodesic $t\mapsto\gamma(t)$, $t\in(0,1]$, can be extended to a geodesic on $t\in(-\infty,1]$. We say that $(M,F)$ is complete if it is both forward and backward complete. According to [2, 35], if $(M,F)$ is either forward or backward complete, then for every two points $x_{1},x_{2}\in M$, there exists a minimal geodesic $\gamma$ from $x_{1}$ to $x_{2}$ such that $d_{F}(x_{1},x_{2})=L_{F}(\gamma)$. Moreover, the closure of a forward (resp., backward) metric ball with finite radius is compact if $(M,F)$ is forward (resp., backward) complete.
@@ 2.2. Measure and polar coordinate system
Let $(M,g)$ be a Riemannian manifold. Given an arbitrary smooth positive measure $\operatorname{\mathfrak{m}}$ on $M$, there always exists some $f\in C^{\infty}(M)$ such that $\operatorname{\mathfrak{m}}=e^{-f}\operatorname{vol}_{g}$. Thus, for any $x\in M$ and any $y\in T_{x}M\backslash\{0\}$,
$\tau(x,y)=f(x),\qquad\mathbf{S}(x,y)=g(y,\nabla f|_{x}),$ (2.12)
see Zhao [36]. According to (2.12), $S$-curvature in the weighted Riemannian setting cannot be of constant sign, unless it vanishes; similar situation occurs on a reversible Finsler manifold as well, see Zhao [36, (2.10)]. However, on irreversible Finsler manifolds, the $S$-curvature can be of constant sign, i.e., either positive or negative. Such phenomenon happens on every Funk metric space endowed with the Busemann–Hausdorff measure, when the $S$-curvature is a positive constant and it depends only on the dimension of the manifold, see Shen [26].
Given a point $o\in M$, set $S_{o}M:=\{y\in T_{o}M:F(o,y)=1\}$. Define the cut value $i_{y}$ of $y\in S_{o}M$ and the injectivity $\mathfrak{i}_{o}$ at $o$ as
@@
In particular, if $(M,F)$ is a Cartan–Hadamard manifold (i.e., a simply connected, forward complete Finsler manifold with $\mathbf{K}\leq 0$), then $\mathfrak{i}_{x}=+\infty$ for every $x\in M$.
According to Zhao and Shen [37, Section 3], given a point $o\in M$, one can define the polar coordinate system $(r,y)$ around $o$, where $r(x):=d_{F}(o,x)$ and $y\in S_{o}M$. In particular, for any $y\in S_{o}M$ and $r\in(0,i_{y})$,
@@
Moreover, it follows by Shen [27, Lemma 3.2.3] that the eikonal equations hold, i.e., $F(\nabla r)=F^{*}({\operatorname{d}}r)=1$ for $\operatorname{\mathfrak{m}}$-a.e. in $M$.
By the polar coordinate system $(r,y)$, the measure $\operatorname{\mathfrak{m}}$ can be disintegrated into the form
@@
where ${\operatorname{d}}\nu_{o}(y)$ is the Riemannian volume form of $S_{o}M$ induced by $g_{y}$. In particular, [37, Lemma 3.1] yields
@@
The following notations are useful in the sequel:
$\mathfrak{s}_{k}(t):=\left\{\begin{array}[]{lll}\frac{\sin(\sqrt{k}t)}{\sqrt{k}},&&\text{ if }k>0,\\ \ \ \ \ t,&&\text{ if }k=0,\\ \frac{\sinh(\sqrt{-k}t)}{\sqrt{-k}},&&\text{ if }k<0,\end{array}\right.\qquad\mathfrak{c}_{k}(t):=\frac{\operatorname{d}}{{\operatorname{d}}t}\mathfrak{s}_{k}(t)=\left\{\begin{array}[]{lll}{\cos(\sqrt{k}t)},&&\text{ if }k>0,\\ \ \ \ \ \ 1,&&\text{ if }k=0,\\ {\cosh(\sqrt{-k}t)},&&\text{ if }k<0.\end{array}\right.$
In view of (2.14), we recall a volume comparison result, see Zhao and Shen [37, Theorem 3.4].
@@
(see [37]) Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}\geq(n-1)k$. For every $o\in M$, there holds
@@
where $(r,y)$ is the polar coordinate system at $o$. In particular, we have
@@
for any $R>0$, with equality for some $R_{0}>0$ if and only if for any $y\in S_{o}M$,
@@
$\hat{\sigma}_{o}(r,y)\leq e^{-\tau({\gamma}_{y}^{\prime}(r))}\mathfrak{s}_{-k^{2}}^{n-1}(r)=e^{-\tau(y)-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r),\quad r\in(0,i_{y}).$ (2.17)
As in Huang, Kristály and Zhao [10], we recall the integral of distortion
@@
where $\mathbb{S}^{n-1}$ is the $(n-1)$-dimensional Euclidean unit sphere.
@@
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $o$ be a fixed point in $M$. Set $r(x):=d_{F}(o,x)$ and $\mu:=r^{-p}\operatorname{\mathfrak{m}}$ for some $p\in\mathbb{R}$. Thus,
@@ 3. Sobolev spaces over Finsler manifolds and approximation
Since $W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ is not a vector space for some $p\in[1,\infty)$, there exists $u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ with $\|-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}=+\infty$, i.e., $\int_{M}F^{*p}(-{\operatorname{d}}u)\operatorname{d\mathfrak{m}}=+\infty$. Suppose by contradiction that $\lambda_{F}(M)<+\infty$. Thus (2.5) yields $\lambda_{F}(M)=\lambda_{F^{*}}(M)$ and hence,
@@
Given any $u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$, there exists a sequence $(u_{k})\subset C^{\infty}_{0}(M)$ such that $\|u-u_{k}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0$. Thus, Hölder’s inequality yields $\|u-u_{k}\|_{W^{1,q}_{\operatorname{\mathfrak{m}}}}\leq\|u-u_{k}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0$, i.e., $u\in W^{1,q}_{0}(M,\operatorname{\mathfrak{m}})$. Moreover, (3.2) combined with Hölder’s inequality again gives
@@
which implies $-u\in W^{1,q}_{0}(M,\operatorname{\mathfrak{m}})$. ∎
@@
The function $f$ is also called $C$-Lipschitz. The minimal $C$ satisfying the above inequality is called the dilatation of $f$, denoted by $\mathsf{dil}(f)$.
@@
For a forward complete Finsler manifold $(M,F)$, one has $C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$.
@@
where $\mathscr{K}\subset M$ is a compact set and $C>0$ is a constant. In particular, if $u$ is nonnegative, so are $u_{l}$’s.
@@
where $\|\cdot\|_{W^{1,p}(\mathbb{B}^{n})}$ is the standard Sobolev norm on $\mathbb{B}^{n}$, see (3.1). In fact, let $h_{\epsilon}$ be the convolution of $h$ with a standard mollifier $\phi_{\epsilon}$, i.e., $h_{\epsilon}:=h*\phi_{\epsilon}$. Note that
@@
Step 2. Since $\operatorname{supp}u$ is compact, there exists a finite coordinate covering $\{(V_{i},\varphi_{i})\}_{i=1}^{N<\infty}$ of $\operatorname{supp}u$ such that $V_{i}\Subset M$ and $\varphi_{i}(V_{i})=\mathbb{B}^{n}$. We may also assume that $\mathscr{K}:=\cup_{i=1}^{N}\overline{V_{i}}$ is a compact subset in $M$. Let $\{\eta_{i}\}_{i=1}^{N}$ be a smooth partition of unity subordinated to $\{V_{i}\}_{i=1}^{N}$. An argument similar to Bao, Chern and Shen [2, (6.2.3)] yields a constant $L\geq 1$ such that for all $i\in\{1,\ldots,N\}$,
$L^{-1}\leq\frac{F^{*}(\varphi^{*}\xi)}{|\xi|_{i}}\leq L,\ \forall\xi\in T^{*}\mathbb{B}^{n}\backslash\{0\},$ (3.3)
@@
In particular, by the uniform convergence we may assume $|h_{i,\epsilon}|\leq C_{1}$, by choosing a larger $C_{1}$ if necessary.
@@
which is a $NC_{1}L$-Lipschitz function with $|u_{\epsilon}|\leq NC_{1}$ and $\operatorname{supp}u_{\epsilon}\subset\cup_{i=1}^{N}V_{i}\subset\mathscr{K}$. Moreover, by a direct calculation one has
@@
In particular, if $u\geq 0$, then $h_{i}\geq 0$ and hence, $h_{i,\epsilon}\geq 0$ due to Step 1. Therefore, $u_{\epsilon}\geq 0$. We are done by letting $\epsilon=1/l$ and $C:=NC_{1}L$. ∎
@@
where $r(x):=d_{F}(o,x)$. Let $u_{\alpha}(x):=-e^{-\alpha r(x)}$ for $\alpha>0$. Then the following statements hold::::
@@
Case 1. Suppose $k=0$, in which case $\mathfrak{s}_{-k^{2}}(r)=r$ and $h\geq 0$. Thus (4.2) combined with (2.14) yields
@@
Thus, for any $p\in[1,n)$, an argument similar to (4.3) yields

### v2

@@ 2.1. Elements from Finsler geometry
In this section we recall some definitions and properties from Finsler geometry; for details see e.g. Bao–Chern–Shen [2] and Shen [27, 28].
@@ 2.1.1. Finsler manifolds
A Minkowski norm on an nn($\geq 2$)-dimensional vector space $V$ is a $C^{\infty}$-function $\phi:V\setminus\{0\}\rightarrow[0,+\infty)$ satisfying:
@@
[…]
The reversibility of a set $U\subset M$, following Rademacher [25], is defined by
@@
Observe that $\lambda_{F}(M)\geq 1,$ with equality if and only if $F$ is reversible, i.e., F(x,y)=F(x,y)F(x,y)=F(x,-y) for all $(x,y)\in TM$. Moreover, $x\mapsto\lambda_{F}(x)$ is continuous on $M$, and hence $\lambda_{F}(M)$ is finite whenever $M$ is compact or $F$ is a Minkowski metric.
@@
where $g_{y}(y,\cdot):=g_{ij}(x,y)y^{i}{\operatorname{d}}x^{j}$. Given a smooth function $u:M\rightarrow\mathbb{R}$, its gradient $\nabla u(x)$ is defined as
@@
This leads to the identity $\langle X,{\operatorname{d}}u\rangle=X(u)=g_{\nabla u}(\nabla u,X)$ valid for any smooth vector field $X$. Here $\langle y,\xi\rangle:=\xi(y)$ denotes the canonical pairing between $T_{x}M$ and $T^{*}_{x}M$.
@@
[…]
with equality if and only if $\xi=\alpha\mathfrak{L}(x,y)$ for some $\alpha>0$. In addition, $F(x,y)=F^{*}(\mathfrak{L}(x,y))$ for any $y\in T_{x}M$. Furthermore, the reversibilities of $F$ and $F^{*}$ coincide, see Huang–Kristály–Zhao [10, Lemma 2.1], i.e.,
@@
The distance function $d_{F}:M\times M\rightarrow[0,\infty)$ is defined as $d_{F}(x_{1},x_{2}):=\inf L_{F}(c),$ where the infimum is taken over all locally Lipschitz curves $c:[0,1]\rightarrow M$ with $c(0)=x_{1}$ and $c(1)=x_{2}$. For any $x_{1},x_{2},x_{3}\in M$, it satisfies $d_{F}(x_{1},x_{2})\geq 0,$ with equality if and only if $x_{1}=x_{2}$, and $d_{F}(x_{1},x_{2})\leq d_{F}(x_{1},x_{3})+d_{F}(x_{3},x_{2})$. Usually $d_{F}(x_{1},x_{2})\neq d_{F}(x_{2},x_{1}),$ unless $F$ is reversible. In fact, one has for every $x_{1},x_{2}\in M$ that
@@
A Finsler manifold $(M,F)$ is called forward complete if every geodesic $t\mapsto\gamma(t)$, $t\in[0,1)$, can be extended to a geodesic defined on $t\in[0,+\infty)$; similarly, $(M,F)$ is called backward complete if every geodesic $t\mapsto\gamma(t)$, $t\in(0,1]$, can be extended to a geodesic on $t\in(-\infty,1]$. We say that $(M,F)$ is complete if it is both forward and backward complete. According to [2], if $(M,F)$ is either forward or backward complete, then for every two points $x_{1},x_{2}\in M$, there exists a minimal geodesic $\gamma$ from $x_{1}$ to $x_{2}$ such that $d_{F}(x_{1},x_{2})=L_{F}(\gamma)$. Moreover, the closure of a forward (resp., backward) metric ball with finite radius is compact if $(M,F)$ is forward (resp., backward) complete.
@@ 2.2. Measure and polar coordinate system
Given a point $o\in M$, set $S_{o}M:=\{y\in T_{o}M:F(o,y)=1\}$. The cut value $i_{y}$ of a direction $y\in S_{o}M$ and the injectivity $\mathfrak{i}_{o}$ at $o$ are defined by
@@
Moreover, if $(M,F)$ is a Cartan–Hadamard manifold (i.e., a simply connected, forward complete Finsler manifold with $\mathbf{K}\leq 0$), then $\mathfrak{i}_{x}=+\infty$ for every $x\in M$.
Given a point $o\in M$, one can define the polar coordinate system $(r,y)$ around $o$, where $r(x):=d_{F}(o,x)$ and $y\in S_{o}M$; see Zhao–Shen [37, Section 3] for example. In particular, for any $y\in S_{o}M$ and $r\in(0,i_{y})$,
@@
Moreover, it follows by [28, Lemma 3.2.3] that the eikonal equations hold, i.e., $F(\nabla r)=F^{*}({\operatorname{d}}r)=1$ for $\operatorname{\mathfrak{m}}$-a.e. in $M$.
In the polar coordinate system $(r,y)$, the measure $\operatorname{\mathfrak{m}}$ decomposes as
@@
where ${\operatorname{d}}\nu_{o}(y)$ is the Riemannian volume form of $S_{o}M$ induced by $g_{y}$. Furthermore, it follows from [37, Lemma 3.1] that
@@
The following notations are useful in the sequel: $\pi/\sqrt{k}:=+\infty$ if $k\leq 0$,
[…]
In view of (2.13), we recall a volume comparison result, see Zhao–Shen [37, Theorem 3.4].
@@
(Zhao–Shen [37]) Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}\geq(n-1)k$. For every $o\in M$, there holds
@@
where $(r,y)$ is the polar coordinate system at $o$. Moreover, for any $R\in(0,\pi/\sqrt{k})$,
@@
with equality for some $R_{0}\in(0,\pi/\sqrt{k})$ if and only if for any $y\in S_{o}M$,
@@
$\hat{\sigma}_{o}(r,y)\leq e^{-\tau({\gamma}_{y}^{\prime}(r))}\mathfrak{s}_{k}^{n-1}(r)\leq e^{-\tau(y)-(n-1)hr}\mathfrak{s}_{k}^{n-1}(r),\quad r\in(0,i_{y}).$ (2.16)
As in [10], we recall the integral of distortion
@@
where $\mathbb{S}^{n-1}$ is the (n1)(n-1)-dimensional Euclidean unit sphere.
@@
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $o$ be a fixed point in $M$. Set $r(x):=d_{F}(o,x)$ and $\mu:=r^{-p}\operatorname{\mathfrak{m}}$ for some $p\in\mathbb{R}$. Then
@@ 3. Sobolev spaces over Finsler manifolds and approximation
Since $W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ is not a vector space for some $p\in[1,\infty)$, there exists $u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ with $\|-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}=+\infty$, i.e., $\int_{M}F^{*p}(-{\operatorname{d}}u)\operatorname{d\mathfrak{m}}=+\infty$. Suppose by contradiction that $\lambda_{F}(M)<+\infty$. Then (2.5) yields $\lambda_{F}(M)=\lambda_{F^{*}}(M)$ and hence,
@@
Given any $u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$, there exists a sequence $(u_{k})\subset C^{\infty}_{0}(M)$ such that $\|u-u_{k}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0$. Hölder’s inequality then yields $\|u-u_{k}\|_{W^{1,q}_{\operatorname{\mathfrak{m}}}}\leq\|u-u_{k}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0$, i.e., $u\in W^{1,q}_{0}(M,\operatorname{\mathfrak{m}})$. Moreover, (3.2) combined with Hölder’s inequality again gives
@@
which implies $u\in W^{1,q}_{0}(M,\operatorname{\mathfrak{m}})$. ∎
@@
The function $f$ is also called $C$-Lipschitz. The minimal $C$ satisfying the above inequality is called the Lipschitz constant or dilatation of $f$, denoted by $\mathsf{dil}(f)$.
@@
Let $(M,F)$ be a forward complete Finsler manifold. Then $C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$.
@@
[…]
@@
where $\|\cdot\|_{W^{1,p}(\mathbb{B}^{n})}$ is the standard Sobolev norm on $\mathbb{B}^{n}$, see (3.1). Let $h_{\epsilon}$ be the convolution of $h$ with a standard mollifier $\phi_{\epsilon}$, i.e., $h_{\epsilon}:=h*\phi_{\epsilon}$. Note that
@@
Step 2. Since $\operatorname{supp}u$ is compact, there exists a finite coordinate covering $\{(V_{i},\varphi_{i})\}_{i=1}^{N<\infty}$ of $\operatorname{supp}u$ such that $V_{i}\Subset M$ and $\varphi_{i}(V_{i})=\mathbb{B}^{n}$. We may also assume that $\mathscr{K}:=\cup_{i=1}^{N}\overline{V_{i}}$ is a compact subset in $M$. Let $\{\eta_{i}\}_{i=1}^{N}$ be a smooth partition of unity subordinated to $\{V_{i}\}_{i=1}^{N}$. An argument similar to [2, (6.2.3)] yields a constant $L\geq 1$ such that for all $i\in\{1,\ldots,N\}$,
[…]
@@
Moreover, by the uniform convergence we may assume $|h_{i,\epsilon}|\leq C_{1}$, by choosing a larger $C_{1}$ if necessary.
@@
which is a $NC_{1}L$-Lipschitz function with $|u_{\epsilon}|\leq NC_{1}$ and $\operatorname{supp}u_{\epsilon}\subset\cup_{i=1}^{N}V_{i}\subset\mathscr{K}$. Moreover, a direct calculation gives
@@
Furthermore, if $u\geq 0$, then $h_{i}\geq 0$ and hence, $h_{i,\epsilon}\geq 0$ due to Step 1. Therefore, $u_{\epsilon}\geq 0$. We are done by letting $\epsilon=1/l$ and $C:=NC_{1}L$. ∎
@@
where $r(x):=d_{F}(o,x)$. Let $u_{\alpha}(x):=-e^{-\alpha r(x)}$ for $\alpha>0$. Then the following statements hold::
@@
Case 1. Suppose $k=0$, in which case $\mathfrak{s}_{-k^{2}}(r)=r$ and $h\geq 0$. Then (4.2) combined with (2.13) yields
@@
Hence, for any $p\in[1,n)$, an argument similar to (4.3) yields